OpenStudy (anonymous):

graph using transformations of the known, basic functions: g(x) = |X-2|^2 + 1

OpenStudy (aum):

Graph y = |x|^2 which would be the same as y = x^2 because y >= 0 for all x. g(x) = |x - 2|^2 + 1 Move the graph of y = |x|^2 2 units to the right and 1 unit up and you will get g(x) = |x - 2|^2 + 1 Replacing x by (x-2) move the graph 2 units to the right. Adding the 1 moves it 1 unit up.

OpenStudy (anonymous):

thank you

OpenStudy (aum):

you are welcome.